# Token Reduction Should Go Beyond Efficiency in Generative Models – From Vision, Language to Multimodality

### v2

@@ 5.5 Towards Dense Prediction Tasks for Vision
Existing works primarily concentrate on compressing the backbone of models to ensure their generalization ability, and few works explore recovering all tokens for dense prediction tasks bolya2023token ; liang2022expediting . It is necessary to develop custom token reduction methods for various downstream dense prediction applications like autonomous driving and robotic control with specific settings and requirements. Lacking these specialized designs would lead to a mismatch and performance drop when deployed in real-world settings. For example, autonomous driving zhang2024sparsead would require displacement and velocity based on occupancy prediction, and robotic control wang2024sparse would demand rotation angle according to the grid map. Therefore, how to develop fast and specialized token reduction strategies tailored for downstream dense prediction tasks is crucial for deployment in practical scenarios.
@@ 5.6 Towards Long Video Applications
Exploiting long videos holds great potential, as processing hours of footage is significantly more labor-intensive and time-consuming than working with short clips. Due to the inherent complexity and resource demands, most current research on long video learning focuses on discriminative tasks such as video understanding lee2024video ; ren2023testa . In contrast, broader applications including long video editing zhang2025adaflow , long video-text retrieval, and narrative-level generation remain largely underexplored. Progress in these areas could have a significant impact on scene editing in video clips, character rendering in movies, and retrieving useful information from numerous videos.
@@
Moreover, token reduction offers a path toward interpretability and efficiency in long video processing jiang2025token . This mimics the human visual system, which does not attend to every frame in detail but instead focuses on salient spatiotemporal changes, such as actions or object movement, while filtering out static, redundant content like backgrounds and stationary objects. Future models should similarly prioritize informative frames and temporal segments, allowing them to reason over extended video sequences with greater efficiency and interpretability.
@@ 5.7 Towards AI for Broader ML and Scientific Domains
Building Biomedical Tokenizers. Recent works exemplify the transformative potential of advanced tokenization methods in the biomedical domain, including protein yuan2025protein ; gao2025foldtoken ; suyunu2025evobpe , genomic eapen2025genomic , and chemical structure yan2024invariant tokenizers. Collectively, these methods illustrate how informed reduction and condensation of input tokens can lead to more effective and interpretable scientific models. For example, traditional tokenizers in EHR foundation models typically treat medical codes as isolated textual units, neglecting their inherent structured and relational context, such as hierarchical relationships, disease co-occurrences, and drug-treatment associations found within biomedical ontologies. To solve this issue, MedTok su2025multimodal integrates textual descriptions and graph-based relational data into a unified tokenization framework. It first uses a language model encoder to extract embeddings from medical code descriptions and employs a graph encoder to capture relational structures from biomedical ontologies. These embeddings are combined into a compact token space through vector quantization, preserving both modality-specific and cross-modality information.
@@
To enhance informativeness and reduce redundancy, MedTok employs a token packing mechanism. It optimizes shared tokens and modality-specific tokens, ensuring that the final tokens encode both shared semantic meaning and modality-specific structure. This process drastically reduces effective vocabulary size, addressing the scalability challenge of 600,000+ medical codes by collapsing redundant representations while preserving critical clinical context. Inspired by adaptive tokenization methods for vision duggal2024adaptive ; kang2023tictok ; yan2024elastictok , future EHR tokenization would be adaptive, enabling the dynamic representation of patients’ medical histories, where the length of the token series for each patient’s history would be directly correlated with the length and complexity. Such adaptive tokenization can significantly improve training and inference efficiency across diverse healthcare systems.
@@
Time-Series Data and Clinical Reasoning. Temporal dynamics form an essential component of clinical reasoning, particularly through longitudinal patient data like lab tests and vital signs. However, current large language models struggle to effectively incorporate time-series inputs due to challenges in temporal tokenization spathis2024first ; anjum2024lipcot ; fang2025tsla ; spathis2024first ; masserano2024enhancing . Future tokenization methods should not only dynamically adjust the number of tokens according to temporal complexity but also selectively focus on time segments most relevant to the clinical context, prompt, or task at hand talukder2024totem . This could enhance training effectiveness and inference accuracy, helping create the next generation of EHR foundation models, which are flexible not only over different tasks or prompts, but also over different data sources, patients, and populations. The complexity and richness of EHR data offer opportunities for AI-driven advancements in patient health outcomes. Future EHR models should support comprehensive reasoning capabilities, encompassing complete patient histories, such as vitals, lab results, diagnoses, and procedures over time. They could facilitate timely disease predictions, accurately forecast chronic disease trajectories, and anticipate patient responses to treatments.
5.8 Algorithm-Hardware Co-Design
While algorithmic advancements in token reduction have achieved impressive computational savings, the next crucial step is to integrate these techniques with hardware-aware design principles. We posit that algorithm-hardware co-design is essential for holistic optimization across the compute stack, considering the interplay between algorithmic choices, hardware architectures (specialized data paths, memory hierarchies, communication fabrics, control logic, etc.), and compiler/runtime support (efficient sparse mapping, dynamic scheduling, irregular-data management, etc.) dong2023heatvit ; parikh2024accelerating .
Currently, co-design efforts targeted at token reduction lag significantly behind pure algorithmic research. This gap is problematic because hardware design needs to balance PPA (power, performance, and area), platform specifics, data movement costs, control overhead, and scalability/reusability 10.1109/ASP-DAC58780.2024.10473968 . Algorithms developed in isolation often generate sparse or irregular compute patterns that general-purpose hardware cannot exploit effectively. Therefore, future research should aim to: 1) Design parameterizable, reconfigurable accelerator modules-such as on-the-fly importance-scoring units and sparse-data pipelines-that natively support token-reduced Transformers. 2) Explore Processing-in-Memory (PIM) architectures to alleviate severe memory bottlenecks caused by dynamic token pruning. By executing scoring operations or partial attention mechanisms within or near memory arrays, PIM can drastically reduce data movement costs and improve end-to-end efficiency.
@@ 6 Conclusion
In this position paper, we have argued that token reduction must evolve beyond a mere efficiency optimization to become a core design principle in generative modeling. We have shown how principled token reduction can address key challenges such as enhancing semantic fidelity in vision-language alignment, curbing verbose reasoning trajectories, preserving long-range coherence, and stabilizing learning dynamics. Looking ahead, we outlined a roadmap of promising directions. We anticipate that embracing token reduction as a holistic, task-aware mechanism will yield the next generation of generative architectures that are not only more efficient but also more robust, interpretable, and aligned with real-world demands.

### v3

@@ 5.5 Towards Dense Prediction Tasks for Vision
Existing works primarily concentrate on compressing the backbone of models to ensure their generalization ability, and few works explore recovering all tokens for dense prediction tasks Bolya and Hoffman (2023); Liang et al. (2022a). It is necessary to develop custom token reduction methods for various downstream dense prediction applications like autonomous driving and robotic control with specific settings and requirements Cao et al. (2025); Pei et al. (2025). Lacking these specialized designs would lead to a mismatch and performance drop when deployed in real-world settings. For example, autonomous driving Zhang et al. (2024a) would require displacement and velocity based on occupancy prediction, and robotic control Wang et al. (2024c) would demand rotation angle according to the grid map. Therefore, how to develop fast and specialized token reduction strategies tailored for downstream dense prediction tasks is crucial for deployment in practical scenarios.
@@ 5.6 Towards Long Video Applications
Exploiting long videos holds great potential, as processing hours of footage is significantly more labor-intensive and time-consuming than working with short clips. Due to the inherent complexity and resource demands, most current research on long video learning focuses on discriminative tasks such as video understanding Lee et al. (2024); Ren et al. (2023). In contrast, broader applications including long video editing Zhang et al. (2025f), long video-text retrieval, and narrative-level generation remain largely underexplored. Progress in these areas could have a significant impact on scene editing in video clips, character rendering in movies, and retrieving useful information from numerous videos.
@@
Moreover, token reduction offers a path toward interpretability and efficiency in long video processing Jiang et al. (2025a). This mimics the human visual system, which does not attend to every frame in detail but instead focuses on salient spatiotemporal changes, such as actions or object movement, while filtering out static, redundant content like backgrounds and stationary objects. Future models should similarly prioritize informative frames and temporal segments, allowing them to reason over extended video sequences with greater efficiency and interpretability.
5.7 Algorithm-Hardware Co-Design
While algorithmic advancements in token reduction have achieved impressive computational savings, the next crucial step is to integrate these techniques with hardware-aware design principles. We posit that algorithm-hardware co-design is essential for holistic optimization across the compute stack, considering the interplay between algorithmic choices, hardware architectures (specialized data paths, memory hierarchies, communication fabrics, control logic, etc.), and compiler/runtime support (efficient sparse mapping, dynamic scheduling, irregular-data management, etc.) Dong et al. (2023); Parikh et al. (2024).
Currently, co-design efforts targeted at token reduction lag significantly behind pure algorithmic research. This gap is problematic because hardware design needs to balance PPA (power, performance, and area), platform specifics, data movement costs, control overhead, and scalability/reusability Pan et al. (2024). Algorithms developed in isolation often generate sparse or irregular compute patterns that general-purpose hardware cannot exploit effectively. Therefore, future research should aim to: 1) Design parameterizable, reconfigurable accelerator modules-such as on-the-fly importance-scoring units and sparse-data pipelines-that natively support token-reduced Transformers. 2) Explore Processing-in-Memory (PIM) architectures to alleviate severe memory bottlenecks caused by dynamic token pruning. By executing scoring operations or partial attention mechanisms within or near memory arrays, PIM can drastically reduce data movement costs and improve end-to-end efficiency.
5.8 Towards Efficient Agentic Systems
Dynamic Memory & Context Engineering. AI agents face a severe context bottleneck where accumulating interaction history not only incurs quadratic computational costs but also degrades reasoning through "lost-in-the-middle" phenomena. Transitioning to active context management is essential Anthropic (2025); systems must distinguish between immutable instructions and transient episodic data. ACON Kang et al. (2025) demonstrate that semantic memory compression can reduce peak token usage while maintaining long-horizon performance. Complementary strategies include hierarchical memory abstraction, which offloads older history to retrieval-based storage, and hybrid masking techniques JetBrains Research (2025). These approaches treat tokens as a finite resource to be optimized on-the-fly, dynamically adjusting the sparsity of the context window based on the complexity of the current reasoning step.
Observation & Tool Pruning. Agents must often process verbose tool outputs (e.g., massive JSON/HTML). Feeding raw data is inefficient; future research should focus on token-aware interaction, where lightweight scorers filter observations to retain only task-relevant features. This preserves critical context bandwidth without sacrificing trajectory accuracy, as evidenced in recent programming agents Xiao et al. (2025b). Furthermore, adaptive truncation strategies that prioritize error traces over standard success logs can significantly improve debugging capabilities while minimizing token consumption.
Communication-Efficient Multi-Agent Systems. At the system level, the aggregate token cost scales with the number of interacting nodes. Enforcing strict token budgets on inter-agent exchange compels information-dense communication. This sparse protocol enhances scalability by reducing redundant chatter and mitigating hallucination loops in collaborative workflows Zhang et al. (2024b); Zou et al. (2025b).
@@ 5.9 Towards AI for Broader ML and Scientific Domains
Building Biomedical Tokenizers. Recent works exemplify the transformative potential of advanced tokenization methods in the biomedical domain, including protein Gao et al. (2025b); Suyunu et al. (2025); Yuan et al. (2025), genomic Eapen (2025), and chemical structure Yan et al. (2024) tokenizers. Collectively, these methods illustrate how informed reduction and condensation of input tokens can lead to more effective and interpretable scientific models. For example, traditional tokenizers in EHR foundation models typically treat medical codes as isolated textual units, neglecting their inherent structured and relational context, such as hierarchical relationships, disease co-occurrences, and drug-treatment associations found within biomedical ontologies. To solve this issue, MedTok Su et al. (2025b) integrates textual descriptions and graph-based relational data into a unified tokenization framework. It first uses a language model encoder to extract embeddings from medical code descriptions and employs a graph encoder to capture relational structures from biomedical ontologies. These embeddings are combined into a compact token space through vector quantization, preserving both modality-specific and cross-modality information.
@@
To enhance informativeness and reduce redundancy, MedTok employs a token packing mechanism. It optimizes shared tokens and modality-specific tokens, ensuring that the final tokens encode both shared semantic meaning and modality-specific structure. This process drastically reduces effective vocabulary size, addressing the scalability challenge of 600,000+ medical codes by collapsing redundant representations while preserving critical clinical context. Inspired by adaptive tokenization methods for vision Duggal et al. (2024); Kang and Lee (2023); Yan et al. (2025), future EHR tokenization would be adaptive, enabling the dynamic representation of patients’ medical histories, where the length of the token series for each patient’s history would be directly correlated with the length and complexity. Such adaptive tokenization can significantly improve training and inference efficiency across diverse healthcare systems.
@@
Time-Series Data and Clinical Reasoning. Temporal dynamics form an essential component of clinical reasoning, particularly through longitudinal patient data like lab tests and vital signs. However, current large language models struggle to effectively incorporate time-series inputs due to challenges in temporal tokenization Spathis and Kawsar (2024); Anjum (2024); Fang et al. (2025b); Spathis and Kawsar (2024); Masserano et al. (2024). Future tokenization methods should not only dynamically adjust the number of tokens according to temporal complexity but also selectively focus on time segments most relevant to the clinical context, prompt, or task at hand Talukder et al. (2024). This could enhance training effectiveness and inference accuracy, helping create the next generation of EHR foundation models, which are flexible not only over different tasks or prompts, but also over different data sources, patients, and populations. The complexity and richness of EHR data offer opportunities for AI-driven advancements in patient health outcomes. Future EHR models should support comprehensive reasoning capabilities, encompassing complete patient histories, such as vitals, lab results, diagnoses, and procedures over time. They could facilitate timely disease predictions, accurately forecast chronic disease trajectories, and anticipate patient responses to treatments.
@@ 6 Conclusion
In this position paper, we have argued that token reduction must evolve beyond a mere efficiency optimization to become a core design principle in generative modeling. We have shown how principled token reduction can address key challenges such as enhancing semantic fidelity in vision-language alignment, curbing verbose reasoning trajectories, preserving long-range coherence, and stabilizing learning dynamics. Looking forward, the roadmap we outlined points to a broad landscape of opportunities, ranging from algorithmic innovations and hardware-algorithm co-design to specialized applications in scientific domains. We anticipate that future work will increasingly focus on constructive compression and reinforcement learning-guided selection, enabling models to autonomously optimize their information bandwidth. Ultimately, by treating token reduction as a holistic and task-aware mechanism, the community can develop next-generation systems that effectively balance scalability with effectiveness, interpretability, and performance.
7 Limitations
While token reduction offers significant benefits, our review identifies critical limitations and trade-offs that must be considered to avoid indiscriminate application.
Information Loss in Dense Prediction
Token reduction methods, particularly pruning, inherently discard information. While this is acceptable for semantic classification or generation, it poses severe risks for dense prediction tasks (e.g., segmentation, object detection) or medical analysis where fine-grained spatial details are crucial. Merging strategies like ToMe Bolya and Hoffman (2023) mitigate this better than pruning, but artifacts often remain at high compression ratios. In scenarios requiring pixel-perfect reconstruction, the trade-off between reduction and precision often favors preserving the full token set. Furthermore, the lack of specialized designs for token recovery often leads to performance mismatches in real-world applications like autonomous driving and robotic control. Future research must therefore explore custom reconstruction mechanisms to ensure these methods can meet the rigorous demands of dense tasks.
Overhead vs. Gain
Dynamic token reduction introduces computational overhead (e.g., scoring networks, predictors). For short sequences or small batch sizes, the cost of computing token importance may outweigh the savings from processing fewer tokens. Furthermore, unstructured pruning can lead to irregular memory access patterns that are inefficient on standard hardware (GPUs/TPUs), potentially negating theoretical FLOPs reductions.
Alternatives: Reduction vs. Retrieval
Critics may argue that techniques like Retrieval-Augmented Generation (RAG) or simply scaling context windows render token reduction unnecessary. However, we argue these are complementary. While RAG selects documents, token reduction operates at the sub-document level, filtering noise within relevant chunks. Similarly, while larger context windows allow for more input, they exacerbate the "lost-in-the-middle" phenomenon; token reduction acts as an attention-sharpening mechanism that helps models utilize these long contexts more effectively.
Appendix A Theoretical Formulation
In this section, we provide a unified mathematical framework for token reduction methods. We formalize the token reduction process into two distinct phases: Compression Criteria (how to evaluate tokens) and Compression Strategies (how to reduce tokens).
A.1 Problem Definition
Given an input sequence $X=[x_{1},x_{2},\dots,x_{N}]\in\mathbb{R}^{N\times D}$, where $N$ represents the sequence length and $D$ denotes the feature dimension, the objective of token reduction is to generate a compressed sequence $X^{\prime}=[x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{M}]\in\mathbb{R}^{M\times D}$, such that $M\ll N$.
The reduction process can be formally defined as a composite function of a scoring criterion $\mathcal{E}$ and a compression strategy $\mathcal{P}$:
$X^{\prime}=\mathcal{P}(X,\mathcal{E}(X))$ (1)
where $\mathcal{E}(X)$ outputs importance scores, clustering assignments, or gradient sensitivities, and $\mathcal{P}$ executes the dimensionality reduction. Figure 2 shows the token reduction pipeline.
A.2 Compression Criteria
The scoring function $\mathcal{E}:X\to\mathcal{S}$ determines the semantic value or redundancy of each token. We broaden the categorization to include gradient and entropy-based metrics alongside standard parametric/non-parametric approaches.
Attention-based Scoring. Utilizing the inherent sparsity of the self-attention mechanism, the importance of a token $x_{i}$ is quantified by the attention it receives. This can be global (averaged across all heads/tokens) or targeted (attention from the special [CLS] token or specific query tokens). The score $s_{i}$ is typically calculated as:
$s_{i}=\sum_{j\in\mathcal{Q}}\text{Attn}(x_{j},x_{i})$ (2)
where $\mathcal{Q}$ is the set of query tokens (e.g., $\mathcal{Q}=\{x_{\text{CLS}}\}$ for classification tasks or $\mathcal{Q}=\{x_{1\dots N}\}$ for global density).
Similarity-based Scoring. This approach assumes that tokens close in the feature space contain redundant information. The criterion calculates pairwise distances to identify clusters or redundant pairs. For tokens $(x_{i},x_{j})$, the metric is typically cosine similarity:
$\text{Sim}(x_{i},x_{j})=\frac{x_{i}\cdot x_{j}}{\|x_{i}\|_{2}\|x_{j}\|_{2}}$ (3)
High similarity scores ($\text{Sim}>\tau$) trigger merging operations. Advanced methods extend this to density-based clustering (e.g., K-Means or DPC-KNN) to identify representative centroids.
Gradient and Entropy-based Scoring. Beyond static feature analysis, recent methods employ dynamic metrics. Gradient-based criteria measure a token’s contribution to the loss function, retaining tokens with high gradient norms ($\|\nabla_{x_{i}}\mathcal{L}\|$). Entropy-based criteria evaluate the uncertainty of the model’s prediction; tokens with low information density (low entropy) are candidates for pruning in early-exit or fast-forwarding frameworks.
Parametric Scoring. Parametric methods introduce a lightweight auxiliary module (e.g., a predictor network $\mathcal{M}_{\phi}$) to explicitly predict token utility:
$S=\mathcal{M}_{\phi}(X)$ (4)
where $S\in[0,1]^{N}$ represents the keep-probability or saliency score. These predictors are trained via Gumbel-Softmax or reinforcement learning (RL) to maximize downstream accuracy while minimizing token count.
A.3 Compression Strategies
[…]
Figure 2: The token reduction pipeline. We formulate reduction as a composite of Criteria $\mathcal{E}$ (scoring) and Strategy $\mathcal{P}$ (pruning/merging).
Once the relationships or scores are established, the compression strategy $\mathcal{P}$ transforms the sequence. We classify these into four primary mechanisms.
Token Pruning (Hard & Soft). Pruning is a selection process that discards tokens based on the criteria $\mathcal{E}$.
$X^{\prime}=\{x_{i}\mid s_{i}\in\text{TopK}(S)\lor s_{i}>\tau\}$ (5)
While Hard Pruning permanently removes tokens, Soft Pruning (or "packaging") aggregates the discarded tokens into a single summary token to preserve residual information, preventing total information loss.
Token Merging & Clustering. Merging aggregates information from a set of tokens $\mathcal{C}=\{x_{1},\dots,x_{k}\}$ identified as similar. This ranges from bipartite matching (merging pairs) to density-based clustering (merging large groups). The merged token $x^{\prime}_{\text{cluster}}$ is a weighted average:
$x^{\prime}_{\text{cluster}}=\frac{\sum_{x_{j}\in\mathcal{C}}w_{j}x_{j}}{\sum_{x_{j}\in\mathcal{C}}w_{j}}$ (6)
where $w_{j}$ tracks the token’s size (number of constituent patches), ensuring proportional representation of fused features.
Transformation-based Compression. These methods reduce sequence length through structural operations, exploiting spatial (image) or temporal (video) priors. Common techniques include:
  • Pooling/Unshuffle: Non-parametric downsampling: $\text{Pool}:\mathbb{R}^{H\times W}\to\mathbb{R}^{\frac{H}{r}\times\frac{W}{r}}$.
  • Convolution: Strided convolution to abstract local neighborhoods: $X^{\prime}=\text{Conv}_{k\times k}(X,s)$.
Token Distillation (Query-based). Distillation employs a set of learnable latent queries $Q\in\mathbb{R}^{M\times D}$ to extract information from the input $X$ via cross-attention mechanisms (e.g., Perceiver Resampler or Q-Former). This decouples output length $M$ from input length $N$:
$X^{\prime}=\text{Softmax}\left(\frac{Q(XW_{K})^{T}}{\sqrt{D}}\right)(XW_{V})$ (7)
This strategy is particularly effective for cross-modal alignment, compressing dense visual features into sparse text-aligned tokens.
[…]
Figure 3: Visualization of token reduction. (a) Image: Visual tokens are pruned based on saliency, retaining only the most salient patches. (b) Text: Low-information stop words (gray) are removed to form a compressed semantic core.
We summarize the general workflow of training-free token reduction in Algorithm 1. We also show a visualization of token reduction in Figure 3.
Algorithm 1 General Training-Free Token Reduction Workflow
1:Input tokens $X\in\mathbb{R}^{N\times D}$, Target ratio $r$
2:Compressed tokens $X^{\prime}\in\mathbb{R}^{M\times D}$
[…]
4:Phase 1: Criteria Calculation ($\mathcal{E}$)
5:if Attention-based then
6:  $S\leftarrow\text{Agg}(\text{AttentionMap}(X))$ $\triangleright$ Agg: Sum/Avg over heads
7:else if Similarity-based then
8:  $A\leftarrow XX^{T}/(\|X\|\|X\|)$ $\triangleright$ Cosine Similarity Matrix
9:  Partition $X$ into sets $\{\mathcal{C}_{1},\dots,\mathcal{C}_{M}\}$ via clustering on AA
10:end if
11:Phase 2: Strategy Execution ($\mathcal{P}$)
12:if Pruning then
13:  $\text{Indices}\leftarrow\text{TopK}(S,M)$
14:  $X^{\prime}\leftarrow\text{Gather}(X,\text{Indices})$
15:else if Merging then
16:  for $m=1$ to $M$ do
17:   $x^{\prime}_{m}\leftarrow\text{WeightedSum}(\mathcal{C}_{m})$
18:  end for
19:end if
20:return $X^{\prime}$
A.4 Controllable Reasoning via Reinforcement Learning
Controllable Reasoning refers to the ability of language models to dynamically adjust the depth and length of their reasoning processes according to user-specified constraints (e.g., exact or maximum token lengths), thereby enabling a tunable trade-off between reasoning efficiency and accuracy. State-of-the-art approaches Luo et al. (2025); Aggarwal and Welleck (2025) typically employ reinforcement learning frameworks Shao et al. (2024), where a multi-objective reward function is designed to optimize correctness and length compliance jointly. Specifically, the reward function incorporates two key objectives:
  1. 1.
    Correctness reward: awarded if the model’s output matches the ground-truth answer;
  2. 2.
    Length penalty: imposed if the generated sequence length deviates from the target length.
Exact Length Control. It requires the model to generate reasoning sequences whose length exactly matches a user-specified target length:
$R_{\text{exact}}(y,y_{\text{t}},n_{\text{t}})=\mathbb{I}(y=y_{\text{t}})-\alpha\cdot|n_{\text{t}}-n_{y}|,$ (8)
where $y$ is the generated sequence, $y_{\text{t}}$ is the ground truth answer, $n_{\text{t}}$ is the target token length, $n_{y}$ is the actual token length of the generated sequence, $\mathbb{I}(\cdot)$ is the indicator function (1 if correct, 0 otherwise), and $\alpha$ is a penalty weight balancing correctness and length.
Maximum Length Control. It controls the model to generate reasoning sequences no longer than a specified upper limit, encouraging efficient reasoning within a token budget:
$R_{\text{max}}(y,y_{\text{t}},n_{\text{t}})=\mathbb{I}(y=y_{\text{t}})\cdot\text{clip}\left(\alpha\cdot(n_{\text{t}}-n_{y})+\delta,0,1\right),$ (9)
where $\text{clip}(\cdot,0,1)$ clamps reward to the range $[0,1]$ and $\delta$ is an offset term to avoid zero reward (typically set to 0.5).
Length Efficiency Optimization. It encourages the model to shorten reasoning length while maintaining correctness, particularly useful for reducing redundancy in long-reasoning models:
$R_{\text{eff}}(y,y_{\text{t}},x)=\frac{\bar{L}_{\text{ref}}(x)}{L(y)}-1+\lambda\left(A(y,y_{\text{t}})-\bar{A}_{\text{ref}}(x)\right),$ (10)
where $\bar{L}_{\text{ref}}(x)$ is the average length of reference model outputs for problem $x$, $A(y,y_{\text{t}})$ is the accuracy function (1 if correct, 0 otherwise), $\bar{A}_{\text{ref}}(x)$ is the average accuracy of the reference model on problem $x$, and $\lambda$ is an accuracy penalty weight to prevent performance degradation from over-compression.
By adjusting hyperparameters (e.g., $\alpha$ and $\lambda$) in the reward function, one can flexibly control the model’s tendency to prioritize correctness versus length, which not only achieves precise length control but also maintains or even improves model performance while significantly reducing reasoning overhead.